OpenStudy (anonymous):

se DeMoivre's Theorem to write the complex number in trigonometric form. (1 - i√3)^3 A) 8(cosπ + isinπ) B) 2(cos(5π/3) + isin(5π/3)) C) 8(cos(5π/3) + isin(5π/3)) D) 2(cos(2π/3) + isin(2π/3))

OpenStudy (anonymous):

let \[z=r \left( \cos \theta+\iota \sin \theta \right)=1-\iota \sqrt{3}\] \[r \cos \theta=1,r \sin \theta=-\sqrt{3}\] square and add \[r=\sqrt{1^2+\left( -\sqrt{3} \right)^2}=2\] divide \[\tan \theta=\frac{ -\sqrt{3} }{ 1 }=-\sqrt{3}=-\tan \frac{ \pi }{ 6 }=\tan \frac{ -\pi }{ 3 }=\tan \left( 2 \pi-\frac{ \pi }{ 3 }\right)=\tan \frac{ 5 \pi }{ 3 }\] \[\theta=\frac{ 5 \pi }{ 3 }\] now you can complete.